A population consist of 2,4,5,9 and 10, with sample size of 3. Compute the mean and variance of the

sampling distribution of the sample mean.


1
Expert's answer
2022-04-06T09:04:46-0400

The mean of the sampling distribution of the sample means is the mean of the population from which the scores were sampled:

"\mu_{\bar x} =\mu=\\
=\cfrac{2+4+5+9+10} {5} =6."

Population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\\
=\begin{Bmatrix}
 2-6, 4-6, 5-6, 9-6, 10-6
\end{Bmatrix}="

"=\begin{Bmatrix}
-4, -2, - 1, 3, 4
\end{Bmatrix},"

"\sigma^2=(-4)^2\cdot \cfrac{1}{5}+(-2)^2\cdot \cfrac{1}{5}+(-1)^2\cdot \cfrac{1}{5}+\\
+3^2\cdot \cfrac{1}{5}+4^2\cdot \cfrac{1}{5}=9.2."


Variance of the sampling distribution of sample means:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{9.2}{3}=3.07."